# The Impact of Question Framing on the Performance of Automatic Occupation Coding

### v2

@@ Abstract
Occupational data play a vital role in research, official statistics, and policymaking, yet their collection and accurate classification remain a challenge. This study investigates the effects of occupational question wording on data variability and the performance of automatic coding tools. We conducted and replicated a split-ballot survey experiment in Germany using two common occupational question formats: one focusing on “job title” (Berufsbezeichnung) and another on “occupational tasks” (berufliche Tätigkeit). Our analysis reveals that automatic coding tools, such as CASCOT and OccuCoDe, exhibit sensitivity to the form and origin of the data. Specifically, these tools were more efficient when coding responses to the job title question format compared with the occupational task format, suggesting a potential way to improve the respective questions for many German surveys. In a subsequent “detailed tasks and duties” question, providing a guiding example prompted respondents to give longer answers without broadening the range of unique words they used. These findings highlight the importance of harmonising survey questions and of ensuring that automatic coding tools are robust to differences in question wording. We emphasise the need for further research to optimise question design and coding tools for greater accuracy and applicability in occupational data collection.
@@ 1 Introduction
Work is a fundamental aspect of human life and a key driver of the global economy. Despite its centrality, the collection and aggregation of occupational information for use in official statistics, research, and policymaking remain challenging. While asking someone, "What do you do?" might seem straightforward in everyday conversations, responses such as "I am a teacher" lack the precision needed for statistical analysis. To inform employment trends, occupational health or migration patterns, occupational data must be collected in a way that ensures both clarity and consistency. Moreover, this information needs to be accurately coded into nationally or internationally recognized classification systems, such as the International Standard Classification of Occupations (ISCO).
In survey methodology, the phrasing of occupational questions affects the clarity and detail of respondent’s answers. Accurate coding requires respondents to effectively translate their job familiarity into sufficiently specific descriptions. Occupational questions, commonly structured in two parts—an initial open-ended summary followed by questions about primary tasks—, have been central to censuses and surveys for decades (Tijdens, 2014). Research has shown that the phrasing of these questions significantly influences the length and specificity of the responses (Martínez et al., 2017), which in turn affects the accuracy and reliability of occupational coding (Conrad et al., 2016).
Traditionally, occupational responses have been manually coded by trained professionals. Although accurate, it is labour-intensive, time-consuming, and prone to human error and variability between coders (Massing et al., 2019). Recent advances in automatic coding methods, from simple text-matching algorithms to machine learning and large language models, aim to address these issues by improving efficiency (Schierholz and Schonlau, 2021). Despite their promise, these technologies still struggle with noisy, ambiguous, or overly concise responses. Efforts continue to enhance these methods through more sophisticated models, larger training datasets, and respondent-assisted coding processes. However, technical developments often overlook a crucial point: the quality of occupational data—whether for input or training—is fundamentally shaped by how survey questions are phrased.
As noted by the US Census, the wording of the occupational questions impacts the quality of answers and can hinder the quality of manual coding (Martínez et al., 2017). It is plausible that the same influence applies to automatic coding processes. However, the literature on automatic occupational coding has largely neglected the role of question wording and data origin, despite its potential to affect the performance of coding algorithms. Understanding how the framing of occupational questions influences both the richness of the data and its suitability for automatic classification is critical to improving the accuracy of these methods.
@@ 2.1 Occupational Question Framing
Efforts to record occupations in a comparable form began in national censuses of the late 18th – 19th centuries (Whitby, 2020). For instance, the 1850 U.S. Census simply asked for each male person’s “profession, occupation, or trade,” but enumerators relied on a 156-word instruction sheet that grew to more than 2,500 words by 1900 (Ruggles et al., 2024). The International Labour Organization (ILO) raised a concern in 1949: "Since many of the problems […] are due to the vagueness or insufficiency of the information furnished, particular attention should be paid to the formulation of the questions referring to the occupation" (ILO, 1949, p. 91). Forty years later, they reiterated the importance of well-designed and tested occupational questions, recommending the use of two questions — job title and main duties (ILO, 1987). Yet formal guidance on how to phrase those questions did not appear until 2012, and even then without methodological backing (ILO, 2012). Without clear standards, survey designers have adopted a wide array of wording strategies for occupational questions.
From the 1950s until 2024, the U.S. Census and the American Community Survey (ACS) asked between one and three occupational questions, altering the first only once, in 2018, from “What kind of work was he doing?” to “What was this person’s main occupation?” (Ruggles et al., 2024). Germany, by contrast, tried sixteen different formulations across the Census and the Mikrozensus (see Table 1A in the Appendix) (Forschungsdatenzentrum, n.d.; Zensus 2022, 2022; Zensus 2011, 2011; Volkszählung 1987, 1987; Forschungsdatenzentrum, n.d.). A broader review on the 33 studies in the United States and Europe (Tijdens, 2014) showed that precise wording of the first question varied although most asked for an “occupation/job title”. In contrast, German panel studies prefer "occupational task"-oriented wording (Schneider et al., 2022), and the Swiss Household Panel revised its German-language item in 2014 while leaving the French, Italian, and English versions untouched (Tillmann et al., 2021). Even a single multinational Labour Force Survey (LFS) is inconsistent: the 2024 Swedish one asked "Into which occupation would you classify your work?", whereas LFS Luxembourg in 2021 - "Which occupational task did you perform during the reporting week? Please state the exact title," and LFS Portugal in 2021 -"What is your occupation? Please be as comprehensive and detailed as possible and describe the functions or tasks you perform" (GESIS - Leibniz Institute for the Social Sciences, 2024).
@@ 2.2 Question framing and data quality
Although most respondents understand a question about their “main occupation or title,” the wording is challenging for those with multiple jobs, no formal title, or an industry-specific label that masks what they actually do (Velkoff et al., 2014). Early ISCO documentation had already catalogued typical shortcomings, including vague responses such as “worker” or “owner” and inflated titles (e.g., a cashier presenting themselves as an accountant) (ILO, 1949). If left unclarified, such answers reach coders in unusable form. According to Geis and Hoffmeyer-Zlotnik (2000) up to one-quarter of entries cannot be coded to the maximum level of detail, around one-tenth are so generic that coders must guess (Ganzeboom, 2010). In addition coders from different agencies agree on only half of the codes, due to both systematic and random errors (Massing et al., 2019).
@@
There have been both theoretical and practical efforts to improve the quality of occupational data. For example, Geis and Hoffmeyer-Zlotnik (2000), a former leader in Germany’s occupational classification efforts, argued that coders first need to know what a person does, not what the job is called. He therefore proposed a three-step module: (i) main task, (ii) clarifying probe, and only then (iii) job title. This task-first logic was later adopted in the Federal Statistical Office’s standards (Hoffmeyer-Zlotnik et al., 2024), yet most contemporary German surveys still use the conventional two questions, presumably to limit respondent burden. Additional refinements such as larger answer boxes (Martínez et al., 2017), illustrative examples (Massing et al., 2019), and targeted interviewer training (Cantor and Esposito, 1992) have been tested with mixed success.
@@
Current research points to a trade-off when designing occupational questions for successful coding. Because respondents are unsure which aspects of their work matter most, many supply only a job title (Kim et al., 2020; Christoph et al., 2020; Massing et al., 2019; Tijdens, 2014). However Martínez et al. (2017) found that asking “What is [Name’s] occupation?” is expected to provide more specific answers than “What kind of work does [Name] do?”. Guiding examples can encourage respondents to provide more detailed occupational descriptions (Massing et al., 2019; Martínez et al., 2017), yet several studies show that example-prompted responses can also become longer, and harder to classify (Conrad et al., 2016; Massing et al., 2019; Kim et al., 2020). According to Cantor and Esposito (1992), coders find that longer occupational descriptions are more likely to include contradictory information, making manual coding less reliable. However, Martínez et al. (2017) and Massing et al. (2019) argue that the issue is more nuanced, as some respondents provide better answers that are easier to code when offered examples.
@@ 2.3 Automatic Coding
Despite the continued reliance on human coders, automatic solutions now classify occupations in seconds rather than hours. In Martínez et al. (2017), trained coders spent 52–126 minutes on 100 cases; an automatic tool can process the same batch almost instantly. Whereas coders draw on job title, duties, industry, education and other hints, most algorithms ingest only the first question (Wan and et al., 2023), and their accuracy, benchmarked against human codes, show considerable variability.
For instance, CASCOT, developed by Jones and Elias (2004), is one of the most well-known multi-language text-matching algorithms, relying on job title inputs. On Dutch job-title data it coded 86% of answers, yet only 52% matched human coders (Belloni and et al., 2016); on translated Chinese occupational data the match rate was only 22% even though 97% of titles received a code (Wan and et al., 2023). While direct comparisons are unavailable, these results suggest that the same tool can perform very differently depending on the survey and question design, even when both use variations of the "What is your job title?" question. A predecessor of OccuCoDe (a German-language machine learning based occupational coding tool) showed similarly uneven performance when coding data from several German surveys (Schierholz and Schonlau, 2021). Coding accuracy shifted with whether the first question asked for job titles or occupational tasks, implying that wording matters - although the direction is still unclear.
Not all algorithms rely on job titles as input. For example, Russ and et al. (2016) developed and tested an ensemble classifier (SOCcer). Their results showed that predictions based on job titles alone had an accuracy of 43%, which increased to 45% when job titles, task descriptions, and industry information were also included. Similarly, LLM4Jobs (a large language model-based tool), tested on full job-posting texts, likewise outperformed CASCOT, which struggles with long inputs (Nan Li and Bie, 2023).
@@
So far, debates about codability (the likelihood of assigning the correct code) are largely centred on the technical specifics of the algorithms. Yet we still lack empirical evidence that the question formats feeding these algorithms yield occupational data of comparable quality, leaving open whether input design, rather than algorithmic sophistication alone, governs the likelihood of assigning the correct code.
@@ 2.4 Research Gap, Research Questions, Hypotheses
There are several inconsistencies in the existing literature regarding occupational data collection. Although occupational surveys have a long tradition, the literature offers no consensus on the optimal wording of open-ended occupation questions. Prior studies show that small variations in phrasing can alter the specificity of responses and, consequently, the effort required for manual coding. Adding illustrative examples typically prompts respondents to write longer answers, yet these more elaborate texts can further complicate coding. The implications of these design choices for automatic coding, and their potential impact on the lexical richness that such tools might exploit, have not been examined systematically.
Our research addresses these gaps by investigating the following research questions: What is the variability in responses due to different occupational question framings? Do these different framings yield similarly rich information?
To answer these questions, we designed, conducted, and replicated a survey experiment in Germany using two versions of occupational questions commonly found in German surveys. We then tested the responses using two automatic coding tools available for the German language. Based on the collected data, we formulated and tested the following hypotheses:
@@ 3.1 Data
We conducted an online experimental study managed by Forsa and simultaneously replicated it using the panel managed by Bilendi. Both firms are commercial survey-research companies that maintain respondent panels. Forsa operates a probability-based panel, recruiting respondents via a random-digit dialling probability sample and subsequently inviting them to participate in specific studies (Güllner and Schmitt, 2004; Schnell and Klingwort, 2024) [^1] Forsa confirmed that occupation is not asked as a permanent panel profile variable; while panellists might have been asked about their job in other studies, those data remain study-specific, so no systematic occupation-related priming could influence our experiment.. Thus, while it is expected to suffer less from self-selection bias, we recognize that other biases, such as differential response rates, may occur. Bilendi, by contrast, is an entirely opt-in panel (non-probabilistic). The Forsa experiment was conducted in Germany between December 7 and 13, 2022. The replication study, conducted by Bilendi, took place over a shorter period, from December 7 to 9, 2022. Both surveys were administered via the Unipark platform, ensuring an identical user experience across both panels. Both companies compensated respondents with digital vouchers for their participation.
@@ 3.2 Experimental Design
  • Question 1"Task": "What is the occupational task that you mainly perform?" (German: “Welche berufliche Tätigkeit üben Sie derzeit hauptsächlich aus?”), currently used in several German panel surveys.

### v3

@@ Abstract
Occupational data play a vital role in research, official statistics, and policymaking, yet their collection and accurate classification remain a challenge. This study investigates the effects of occupational question wording on data variability and the performance of automatic coding tools. We conducted and replicated a split-ballot survey experiment in Germany using two common occupational question formats: one focusing on “job title” (Berufsbezeichnung) and another on “berufliche Tätigkeit” (loosely translated as occupation or occupational task). Our analysis reveals that automatic coding tools, such as CASCOT and OccuCoDe, exhibit sensitivity to the form and origin of the data. Specifically, these tools were more efficient when coding responses to the job title question format than the occupational task format, suggesting a potential way to improve the respective questions for many German surveys. In a subsequent “detailed tasks and duties” question, providing a guiding example prompted respondents to give longer answers without broadening the range of unique words they used. These findings highlight the importance of harmonising survey questions and and ensuring that automatic coding tools are robust to differences in question wording. Further research is needed to optimise question design and coding tools for greater accuracy and applicability in occupational data collection.
@@ 1 Introduction
Work is a fundamental aspect of human life and a key driver of the global economy. Despite its centrality, the collection and aggregation of occupational information for use in official statistics, research, and policymaking remain challenging. While asking someone, “What do you do?” might seem straightforward in everyday conversations, responses such as “I am a teacher” lack the precision needed for statistical analysis. To inform employment trends, occupational health or migration patterns, occupational data must be collected in a way that ensures both clarity and consistency. Moreover, this information needs to be accurately coded into nationally or internationally recognized classification systems, such as the International Standard Classification of Occupations (ISCO).
In survey methodology, the phrasing of occupational questions affects the clarity and detail of respondents’ answers. Accurate coding requires respondents to effectively translate their job familiarity into sufficiently specific descriptions. Occupational questions, commonly structured in two parts—an initial open-ended summary followed by questions about primary tasks—, have been central to censuses and surveys for decades (Tijdens2014). Research has shown that the phrasing of these questions significantly influences the length and specificity of the responses (Martinez2017), which in turn affects the accuracy and reliability of occupational coding (conrad2016classifying).
Traditionally, occupational responses have been manually coded by trained professionals. Although accurate, it is labour-intensive, time-consuming, and prone to human error and variability between coders (Massing2019). Recent advances in automatic coding methods, from simple text-matching algorithms to machine learning and large language models, aim to address these issues by improving efficiency (schierholz2021machine). Despite their promise, these technologies still struggle with noisy, ambiguous, or overly concise responses. Efforts continue to enhance these methods through more sophisticated models, larger training datasets, and respondent-assisted coding processes. However, technical developments often overlook a crucial point: the quality of occupational data—whether for input or training—is fundamentally shaped by how survey questions are phrased.
As noted by the US Census, the wording of the occupational questions impacts the quality of answers and can hinder the quality of manual coding (Martinez2017). It is plausible that the same influence applies to automatic coding processes. However, the literature on automatic occupational coding has largely neglected the role of question wording and data origin, despite its potential to affect the performance of coding algorithms. Understanding how the framing of occupational questions influences both the richness of the data and its suitability for automatic classification is critical to improving the accuracy of these methods.
@@ 2.1 Occupational Question Framing
Efforts to record occupations in a comparable form began in national censuses of the late 18th – 19th centuries (whitby2020sum). For instance, the 1850 U.S. Census simply asked for each male person’s “profession, occupation, or trade,” but enumerators relied on a 156-word instruction sheet that grew to more than 2,500 words by 1900 (ipums_usa_v15). The International Labour Organization (ILO) raised a concern in 1949: "Since many of the problems […] are due to the vagueness or insufficiency of the information furnished, particular attention should be paid to the formulation of the questions referring to the occupation" (ilo1949isco, p. 91). Forty years later, they reiterated the importance of well-designed and tested occupational questions, recommending the use of two questions — job title and main duties (ilo1987). Yet formal guidance on how to phrase those questions did not appear until 2012, and even then without methodological backing (ilo2012isco). Without clear standards, survey designers have adopted a wide array of wording strategies for occupational questions.
From the 1950s until 2024, the U.S. Census and the American Community Survey (ACS) asked between one and three occupational questions, altering the first only once, in 2018, from “What kind of work was he doing?” to “What was this person’s main occupation?” (ipums_usa_v15). Germany, by contrast, tried sixteen different formulations across the Census and the Mikrozensus (see Table 1A in the Appendix) (forschungsdatenzentrum_vz1970; forschungsdatenzentrum_mikrozensus; zensus_1987; zensus2011; zensus2022). A broader review on the 33 studies in the United States and Europe (Tijdens2014) showed that precise wording of the first question varied although most asked for an “occupation/job title”. In contrast, German panel studies prefer "occupational task"-oriented wording (Schneider2022), and the Swiss Household Panel revised its German-language item in 2014 while leaving the French, Italian, and English versions untouched (Tillmann2021). Even a single multinational Labour Force Survey (LFS) is inconsistent: the 2024 Swedish one asked "Into which occupation would you classify your work?", whereas LFS Luxembourg in 2021 - "Which occupational task did you perform during the reporting week? Please state the exact title," and LFS Portugal in 2021 -"What is your occupation? Please be as comprehensive and detailed as possible and describe the functions or tasks you perform" (EU-LFS-questionnaires).
@@ 2.2 Question framing and data quality
Although most respondents understand a question about their “main occupation or title,” the wording is challenging for those with multiple jobs, no formal title, or an industry-specific label that masks what they actually do (ACS2014). Early ISCO documentation had already catalogued typical shortcomings, including vague responses such as “worker” or “owner” and inflated titles (e.g., a cashier presenting themselves as an accountant) (ilo1949isco). If left unclarified, such answers reach coders in unusable form. According to geis2000stand, up to one-quarter of entries cannot be coded to the maximum level of detail, around one-tenth are so generic that coders must guess (Ganzeboom2010). In addition coders from different agencies agree on only half of the codes, due to both systematic and random errors (Massing2019).
@@
There have been both theoretical and practical efforts to improve the quality of occupational data. For example, geis2000stand building on pappiSozialstrukturanalysenMitUmfragedaten1979, former leaders in Germany’s occupational classification efforts, argued that coders first need to know what a person does, not what the job is called. He therefore proposed a three-step module for coding into ISCO: (i) main task, (ii) clarifying probe, and only then (iii) job title. This task-first logic was later adopted in the Federal Statistical Office’s standards (demographische2024standards), though asking three questions can be overly time-consuming (Schneider2022). Shortened versions are widely used. One option, recommended by zuell2015, is to merge main task and clarifying probe into a single question before asking for a job title. An alternative is to ask for a job title in a first question before, optionally, requesting information about the job tasks; see Table LABEL:tab:appendix-1B in the Appendix for example wordings from several large German surveys. Additional refinements such as larger answer boxes (Martinez2017), illustrative examples (Massing2019), and targeted interviewer training (cantor1992) have been tested with mixed success.
@@
Current research points to a trade-off when designing occupational questions for successful coding. Because respondents are unsure which aspects of their work matter most, many supply only a job title (KKM20; Christoph2020; Massing2019; Tijdens2014). However Martinez2017 found that asking “What is [Name’s] occupation?” is expected to provide more specific answers than “What kind of work does [Name] do?”. When it comes to the follow-up question detailing tasks and duties, guiding examples can encourage respondents to provide more details (Massing2019; Martinez2017), yet several studies show that example-prompted responses can also become longer, and harder to classify (conrad2016classifying; Massing2019; KKM20). According to cantor1992, coders find that longer occupational duties and tasks descriptions are more likely to include contradictory information, making manual coding less reliable. However, Martinez2017 and Massing2019 argue that the issue is more nuanced, as some respondents provide better answers that are easier to code when offered examples.
@@ 2.3 Automatic Coding
Despite the continued reliance on human coders, automatic solutions have become a routine part of occupational data co-production in both general surveys and official statistics production (sebastian_kocar_occupation_2025). These automatic solutions vary along several dimensions: the type of technology applied, interactive coding during data collection vs. post-survey data processing,fully automated vs. computer-assisted coding systems, and the type of occupational information on which coding is based.
First, early automatic coding relied on dictionary-based approaches (lyberg_automated_1992). These were followed by more complex rule-based systems that allowed for partial matching, incorporated common abbreviations, spelling errors, or accounted for the absence of certain words in occupational descriptions. CASCOT, developed by jones2004cascot, is one of the most well-known multilingual text-matching algorithms of this type. Similar dictionary-based approaches have been used in the German Mikrozensus (bernhard_hochstetter_berufsklassifizierung_nodate), or commercial tools developed by Verian (cataneda_russel_collecting_2024). Advances in computational capacity enabled the development of more complex methods, ranging from ensemble classifiers such as SOCcer (Russ2016), to machine-learning-based methods (schierholz2021machine), transformer models such as DistilKoBERT (kim_occupation_2024), and recently large language models such as SOCBot (sturgisSOCbotUsingLarge2025). These developments have not replaced earlier methods; rather, dictionary- and machine-learning-based approaches remain in active use.
Second, occupational information is frequently collected through surveys, although other sources, such as visa applications or job search platforms, exist. Coding can occur after data collection is completed (for example, when paper census questionnaires are processed post-survey), or during the data collection process itself, as in interactive survey designs (schierholz2018occupation; sebastian_kocar_occupation_2025; sturgisSOCbotUsingLarge2025). Some tools, such as CASCOT or OccuCoDe (Simson2023), can be deployed in both settings.
Third, automatic solutions vary by the degree of automation. Fully automatic systems assign the most likely occupational code without human intervention, whereas semi-automatic systems provide suggestions (or a look-up list) that are reviewed and confirmed by an expert coder or by respondents themselves. Fully automatic solutions include PROCODE (savic_procode_2020) or the LLM-based tool by bach_job_2025, although many semi-automatic tools (such as CASCOT, OccuCoDe, or SOCcer) can also be used in a fully automatic mode.
Finally, automatic coding tools differ by the type of occupational information they are designed to process. Despite advances in computational capacity, many existing systems are optimised for short textual inputs rather than longer descriptions. Short answers (typically responses to questions such as “What is your job title?”) are commonly used to generate coding suggestions, while follow-up questions about tasks and duties are often reserved for office coding (sebastian_kocar_occupation_2025). Tools such as CASCOT, OccuCoDe, Verian, PROCODE, SOCBot, Klassifikationsserver by Mikrozensus, primarily rely on short inputs and may not reliably process longer, more detailed descriptions of tasks and duties. A smaller number of tools, including SWEEP (sebastian_kocar_occupation_2025) and SOCcer, are able to incorporate additional information such as tasks or industry alongside job titles.
@@
So far, debates about automatic codability (the likelihood of assigning the correct code) are largely centred on the technical specifics of the algorithms. Yet we still lack empirical evidence that the question formats feeding these algorithms yield occupational data of comparable quality, leaving open whether input design, rather than algorithmic sophistication alone, governs the likelihood of assigning the correct code.
@@ 2.4 Research Gap, Research Questions, Hypotheses
Against this background, several inconsistencies can be observed in the existing literature on occupational data collection. Although occupational surveys have a long tradition, the literature offers no consensus on the optimal wording of open-ended occupation questions. Prior studies show that small variations in phrasing can alter the specificity of responses and, consequently, the effort required for manual coding. Adding illustrative examples typically prompts respondents to write longer answers, yet these more elaborate texts can further complicate coding. What remains largely unexamined is how these design choices affect the structure of occupational information in ways that are relevant for automatic coding, including their potential implications for the lexical material such tools are able to exploit.
Our research addresses these gaps by investigating the following research questions: What is the variability in responses due to different occupational question framings? Do these different framings yield similarly rich information? Do different question framings produce information that is equally conducive to automatic coding?
To answer these research questions, we designed, conducted, and replicated a survey experiment in Germany using two versions of occupational questions commonly found in German surveys. We then tested the responses using two automatic coding tools available for the German language. A focus is on the starting question because the coding tools we use were not developed to process answers from a sequence of occupational questions. Based on the collected data, we formulated and tested the following hypotheses:
@@ 3.1 Data
We conducted an online experimental study managed by Forsa and simultaneously replicated it using the panel managed by Bilendi. Both firms are commercial survey-research companies that maintain respondent panels. Forsa operates a probability-based panel, recruiting respondents via a random-digit dialling probability sample and subsequently inviting them to participate in specific studies (Guellner2004; Schnell2024) [^1] Forsa confirmed that occupation is not asked as a permanent panel profile variable; while panellists might have been asked about their job in other studies, those data remain study-specific, so no systematic occupation-related priming could influence our experiment.. Thus, while it is expected to suffer less from self-selection bias, we recognize that other biases, such as differential response rates, may occur. Bilendi, by contrast, is an entirely opt-in panel (non-probabilistic). The Forsa experiment was conducted in Germany between December 7 and 13, 2022. The replication study, conducted by Bilendi, took place over a shorter period, from December 7 to 9, 2022. Both surveys were administered via the Unipark platform, ensuring an identical user experience across both panels. Both companies compensated respondents with digital vouchers for their participation.
@@ 3.2 Experimental Design
  • Question 1"Task": "What is the occupational task that you mainly perform?" (German: “Welche berufliche Tätigkeit üben Sie derzeit hauptsächlich aus?”), apart from very minor deviations this question text is generally recommended by the German Federal Statistical Office and other institutes (demographische2024standards). Many large panel surveys adopt the task-focused phrase “berufliche Tätigkeit” in their questions, though typically these surveys include job titles as examples or request a detailed description of the occupational tasks.
    According to geis2000stand, the Question 1"Task" followed by question 2 (“Please describe this occupational task in detail.”) is recommended to collect information about the work (tasks and duties) performed by a respondent. We note, however, that the German question text is rather hard to translate to English. Many respondents may just understand the first question as an unusual or formal way of asking “What is your main occupation?”, not noting the intended minor emphasis on which task they do in their job.